X/a+y/b=a^2+b^2 and x/a^2+y/b^2=a+b then value of x

Grade:8

1 Answers

Arun
5 years ago
Dear Kajal
 
2in x/a + y/b = a + b we get x=a2substituting y=b2(b-a)After cancelling a(b-a) on both sides we get y=b2a(b-a)y=ab2b2-a3)y=ab2............(6)subtracting 5 from 6 we get(ab-a2b2y=2a2x+a2.........(5)b3+ab2b2x+aby=a2........(3)Multiplying equation 3 with b we getb2b+ab2..........(4)bx+ay=a2b2y=2a2x+a2Given Equations arex/a^2 + y/b^2 = 2..........(1)x/a + y/b = a + b...........(2)Simplifying equations (1) and (2) by taking lcm we getb